# Aurora-M: The First Open Source Multilingual Language Model Red-teamed according to the U.S. Executive Order

### v2

@@ Abstract
Pretrained language models underpin several AI applications, but their high computational cost for training limits accessibility. Initiatives such as BLOOM and StarCoder aim to democratize access to pretrained models for collaborative community development. However, such existing models face challenges: limited multilingual capabilities, continual pretraining causing catastrophic forgetting, whereas pretraining from scratch is computationally expensive, and compliance with AI safety and development laws. This paper presents Aurora-M, a 15B parameter multilingual open-source model trained on English, Finnish, Hindi, Japanese, Vietnamese, and code. Continually pretrained from StarCoderPlus on 435 billion additional tokens, Aurora-M surpasses 2 trillion tokens in total training token count. It is the first open-source multilingual model fine-tuned on human-reviewed safety instructions, thus aligning its development not only with conventional red-teaming considerations, but also with the specific concerns articulated in the Biden-Harris Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence. Aurora-M is rigorously evaluated across various tasks and languages, demonstrating robustness against catastrophic forgetting and outperforming alternatives in multilingual settings, particularly in safety evaluations. To promote responsible open-source LLM development, Aurora-M and its variants are released here.
Taishi Nakamura1111  Equal contributionMayank Mishra21Simone Tedeschi3,41Yekun Chai5Jason T StillermanFelix Friedrich6,7Prateek Yadav8Tanmay LaudVu Minh Chien9Terry Yue Zhuo10,11Diganta Misra12,13Ben Bogin14Xuan-Son Vu15,16,17Marzena Karpinska18Arnav Varma DantuluriWojciech KusaTommaso FurlanelloRio Yokota1Niklas MuennighoffSuhas Pai19Tosin Adewumi20Veronika LaippalaXiaozhe Yao21Adalberto JuniorAlpay Ariyak22,23Aleksandr Drozd24Jordan Clive25Kshitij Gupta12Liangyu ChenQi Sun1Ken TsuiNoah PersaudNour FahmyTianlong Chen8Mohit Bansal8Nicolò Monti26Tai Dang18Ziyang Luo27Tien-Tung Bui28Roberto Navigli3Virendra Mehta29Matthew Blumberg30222  Equal mentoringVictor May31,322Huu Nguyen322Sampo Pyysalo332
1Tokyo Institute of Technology  2MIT-IBM Watson Lab  3Sapienza University of Rome  4Babelscape  5LAION  6TU Darmstadt  7hessian.AI  8UNC Chapel-Hill  9Detomo Inc.  10CSIRO’s Data61  11Monash University  12Mila - Quebec AI Institute  13Carnegie Mellon University  14Allen Institute for AI  15WASP Media & Language  16Umeå University  17DeepTensor AB  18University of Massachusetts Amherst  19Hudson Labs  20Luleå University of Technology  21ETH Zurich  22RunPod  23OpenChat  24RIKEN CCS  25Chattermill AI  26ASC27  27Hong Kong Baptist University  28DopikAI JSC  29University of Trento  30GridRepublic  31Chegg  32Ontocord.AI  33University of Turku
taishi.nakamura@rio.gsic.titech.ac.jpmayank.mishra2@ibm.comtedeschi@diag.uniroma1.itpraty@cs.unc.edudiganta.misra@mila.quebecmayvic@gmail.comhuu@ontocord.aisampo.pyysalo@utu.fi
Disclaimer: Please be advised that this paper may contain examples of text that could be considered offensive or culturally insensitive. These examples are used only for illustrative and analytical purposes and do not reflect the views of the authors or the publisher.
@@ 1 Introduction
Large Language Models (LLMs) are fundamental tools in artificial intelligence, powering applications such as machine translation, text summarization, dialogue systems, and code generation. These LLMs are pretrained on extensive text data to enhance downstream task-specific adaptation. However, the excessive computational expense of pretraining LLMs creates barriers to access, constraining wider development.
Open-source initiatives such as BLOOM (Scao et al., 2023), StarCoder (Li et al., 2023a), StarCoder-2 (Lozhkov et al., 2024), Pythia (Biderman et al., 2023), and OLMo (Groeneveld et al., 2024; Soldaini et al., 2024) have emerged to democratize access to pretrained LLMs. These initiatives stimulate innovation, allowing researchers and developers to leverage existing advancements. However, despite their contributions, several significant challenges persist in the domain of open-source LLM development.
Primarily, several studies (Bang et al., 2023; Jiao et al., 2023; Hendy et al., 2023; Huang et al., 2023) have underscored the ongoing struggle of LLMs with non-English texts, particularly in low- or extremely low-resource languages. Given that the training data predominantly consists of English, as noted for instance by Brown et al. (2020) who reported that English accounts for 93% of GPT-3’s training corpus, there is a pressing need to promote the development of multilingual models (Chai et al., 2023) to democratize LLMs and alleviate performance disparities across different languages. Secondly, continual pretraining – a technique involving further updating pretrained models on new data distributions to enhance their capabilities – poses a significant challenge. While this approach holds promise for computational saving and performance improvement, it often leads to catastrophic forgetting, where the model loses previously acquired knowledge. This challenge is exacerbated when considering the continual pretraining of models across a diverse array of grammatical and lexical structures. Lastly, ensuring compliance with recent regulations mandating safe and secure AI development practices represents another critical aspect often overlooked in open-source LLM development, specifically, for multilingual models.
This paper presents Aurora-M, a novel open-source multilingual Large Language Model (LLM) with 15 billion parameters, tailored to address the aforementioned limitations. Aurora-M is designed to cater to six linguistically diverse languages: English, Finnish, Hindi, Japanese, Vietnamese, and code. Aurora-M is continually pretrained from the StarCoderPlus model (Li et al., 2023a) on an extensive dataset comprising 435 billion tokens, resulting in a total training token count of an impressive 2 trillion tokens. This rigorous pretraining regimen equips Aurora-M with a comprehensive understanding of diverse languages and code. Moreover, safety is a fundamental design principle of Aurora-M. It stands out as the first open-source multilingual LLM fine-tuned on a comprehensive collection of human-reviewed safety instructions addressing concerns in the Biden-Harris Executive Order on Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence (WhiteHouse, 2023). This fine-tuning process not only addresses conventional red-teaming concerns (Ganguli et al., 2022; Perez et al., 2022; Zhuo et al., 2023; Ge et al., 2023) aimed at testing system vulnerabilities, but also aligns with the specific safety and security guidelines outlined in the Order.
Our contributions can be succinctly summarized as follows.
@@
  • We construct a new red-teaming dataset, named "The Biden-Harris Redteam Dataset," tailored to address concerns outlined in the Executive Order along with typical safety concerns. We then fine-tune Aurora-M on this dataset and evaluate on several safety benchmarks.
@@ 2 Datasets
Data Curation.
The continual pretraining process for training Aurora-M followed a carefully designed two-stage curriculum, as shown in Fig. 2. In the first stage, termed as Continual Auxiliary Pretraining (CAP), a large corpus of general multilingual web data was used to expose the model to diverse data, laying a robust foundation for subsequent training. The second stage, termed as Continual Alignment Tuning (CAT) employed a strategic data-mixing approach to bolster the model’s performance in targeted areas and align it with our predefined objectives. Following Taylor et al. (2022); Li et al. (2023b), we also mixed in publicly available instruction tuning datasets in both stages of training.
In CAP, we incorporated 377B tokens of processed and filtered web data from various sources, including Stack (Kocetkov et al., 2022), RefinedWeb (Penedo et al., 2023), RedPajama (Together, 2023), and a subset of the Pile (Gao et al., 2020). Additionally, multilingual data from HPLT (de Gibert et al., 2024), MC4 (Zhu et al., 2023a), Paracrawl (Ghussin et al., 2023), OSCAR (Abadji et al., 2022), along with Wikipedia (Foundation, ), and instruction tuning data from sources such as OpenAssistant (Köpf et al., 2023), APIBench (Patil et al., 2023), and OIG (LAION, 2023) were included.
For CAT, we opted for a greater percentage of code and a changed mix of high-quality public instruction datasets (Mishra et al., 2022a; Ding et al., 2023; Ivison et al., 2023), encompassing coding (Luo et al., 2023; Mishra et al., 2023a) and mathematical reasoning (Yu et al., 2023; Mishra et al., 2023b). The intention was to not overfit on the high quality instruction data, and thus the high quality data was used in CAT only. We also subsampled data from CAP for quality, as described below. Furthermore, we introduced a new safety instructions dataset named Biden-Harris Redteam, detailed in Section 4. The total dataset size for CAT is 58B tokens. Please refer to Fig. 2 for the distribution of languages in both training stages. The complete list of datasets is available in Appendix B.
@@
To remove toxic content and low-quality text, we applied filters similar to those used in Nguyen et al. (2023b); Scao et al. (2023), such as stop-word proportions and text length. For all web text, we followed a process akin to Penedo et al. (2023) to remove low-quality content, including duplicate headers and footers. Additionally, in the CAT dataset, we further filtered web text with high proportions of symbols and numbers. In the case of RefinedWeb (Penedo et al., 2023), we utilized the RedPajama (Together, 2023) fastText classifier to retain English webpages resembling "high-quality" content similar to Wikipedia-linked articles. We trained and employed a similar classifier to filter other languages in our dataset, except for Finnish, where the procedure caused over-filtering, resulting in an excessively low sample volume post-filtering. To further enhance the quality of the RefinedWeb data, we adopted an approach detailed in Rönnqvist et al. (2021). We trained a fastText classifier [^1] Similar to https://github.com/TurkuNLP/register-labeling?tab=readme-ov-file and selectively subsampled web pages with over-represented registers, aiming to retain more "rare" text (e.g., lyrical or poetic text). This filtering process was specifically applied to English text due to the prohibitive slowness of our multilingual classifiers. Addressing this limitation represents an area for future research.
@@ 3 Model Training
For the training of Aurora-M, we maintained a consistent batch size of 2048 and a sequence length of 2048 tokens. The learning rate was linearly warmed up to $10^{-4}$ over 2,000 steps, followed by a cosine decay scheduler set to decay the learning rate to $10^{-5}$ by 120,000 steps. while optimization utilized the AdamW optimizer (Kingma & Ba, 2017; Loshchilov & Hutter, 2019) with coefficients $\beta_{1}=0.9$ and $\beta_{2}=0.95$. Additionally, Megatron-LM’s distributed optimizer with mixed precision training (Micikevicius et al., 2018) was used. Further training details can be found in the Appendix A.
@@ 4 Safety
  • Attempts to circumvent red-teaming controls.
With these main categories in mind, we curated the Biden-Harris Redteam Dataset comprising 5000 red-teaming instructions, human-reviewed, and edited instruction-response pairs to address lawfulness and safety concerns, including those outlined in the Executive Order (WhiteHouse, 2023). The instructions were sourced both by filtering the human preference dataset on harmlessness from Anthropic (Bai et al., 2022; Ganguli et al., 2022) as well as by means of utilizing semi-automatic template-based methods to generate additional prompts. Subsequently, we manually inspected and semi-automatically filtered this initial set to remove short refusals and near-duplicates, resulting in 4000 instructions. To address potential harmful responses by Aurora-M in the first stage of pretraining, we also used an approximately 1000 instructions subset and hand-wrote or created continuations by this version of Aurora-M. Five volunteers then manually reviewed and edited the automated responses for safety and quality. We utilized the resultant approximately 5000 instructions dataset for instruction-tuning (referred to as the Biden-Harris Redteam Dataset) of Aurora-M and evaluated its safety levels on various safety evaluation datasets both before and after the instruction-tuning step. Details and results are provided in Section 5. Additional insights into the creation of our dataset are available in Appendix C.
@@ 5.1 Evaluation Setup
English Evaluation Datasets.
We used the Language Model Evaluation Harness (Gao et al., 2022). We evaluated question answering tasks, including OpenBookQA (Mihaylov et al., 2018) and TriviaQA (Joshi et al., 2017), natural language inference with HellaSwag (Zellers et al., 2019), machine reading comprehension with SQuAD2.0 (Rajpurkar et al., 2018), XWINO (Tikhonov & Ryabinin, 2021), and arithmetic reasoning with GSM8K (Cobbe et al., 2021) using 8-shot inference.
Japanese Evaluation Datasets.
Following swallow-llama333swallow-llama: https://tokyotech-llm.github.io/swallow-llama, we utilized llm-jp-eval (Han et al., 2024) and the JP Language Model Evaluation Harness [^4] https://github.com/Stability-AI/lm-evaluation-harness. llm-jp-eval utilizes JCommonsenseQA (JCom) (Kurihara et al., 2022) to evaluate multiple choice question answering, JEMHopQA (JEMHop) (Ishii et al., 2023) and NIILC (Sekine, 2003) for free-form question answering, and JSQuAD (Kurihara et al., 2022) for machine reading comprehension using 4-shot inference. JP Language Model Evaluation Harness evaluates automatic summarization on XL-Sum (Hasan et al., 2021) using 1-shot inference, arithmetic reasoning on MGSM (Shi et al., 2023) using 4-shot inference, and Japanese-English and English-Japanese machine translation on WMT 2020 Japanese $\leftrightarrow$ English (Barrault et al., 2020) using 4-shot inference.
Finnish Evaluation Datasets.
Hindi and Vietnamese Evaluation Datasets.
We used the mlmm evaluation666mlmm-evaluation: https://github.com/nlp-uoregon/mlmm-evaluation for evaluation. We evaluated AI2 Reasoning Challenge (Clark et al., 2018), HellaSwag, MMLU (Hendrycks et al., 2021a), and TruthfulQA (Lin et al., 2022) using 0-shot inference. ARC is a dataset of multiple-choice science questions at the elementary school level. HellaSWAG is a dataset for studying grounded commonsense inference. Each question has four choices about what happens next in the scene. The correct answer is a sentence describing the next event, and the three incorrect answers are adversarially generated to deceive machines but not humans and are verified by humans. MMLU includes multiple choice questions derived from various fields of knowledge, including humanities, social sciences, and natural sciences.

### v3

# Aurora-M: Open Source Continual Pre-training for Multilingual Language and Code

@@ Abstract
Pretrained language models are an integral part of AI applications, but their high computational cost for training limits accessibility. Initiatives such as Bloom and StarCoder aim to democratize access to pretrained models for collaborative community development. Despite these efforts, such models encounter challenges such as limited multilingual capabilities, risks of catastrophic forgetting during continual pretraining, and the high costs of training models from scratch, alongside the need to align with AI safety standards and regulatory frameworks. This paper presents Aurora-M, a 15B parameter multilingual open-source model trained on English, Finnish, Hindi, Japanese, Vietnamese, and code. Continually pretrained from StarCoderPlus on 435B additional tokens, Aurora-M surpasses 2T tokens in total training token count. It is the first open-source multilingual model fine-tuned on human-reviewed safety instructions, thus aligning its development not only with conventional red-teaming considerations, but also with the specific concerns articulated in the Biden-Harris Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence. We evaluate Aurora-M across a wide range of tasks and languages, showcasing its robustness against catastrophic forgetting and its superior performance in multilingual settings, particularly in safety evaluations. We open-source Aurora-M and its variants to encourage responsible open-source development of large language models at https://huggingface.co/aurora-m.
Aurora-M: Open Source Continual Pre-training for Multilingual Language and Code
Taishi Nakamura∗1, Mayank Mishra∗2, Simone Tedeschi∗3,4, Yekun Chai5 Jason T Stillerman, Felix Friedrich6,7, Prateek Yadav8, Tanmay Laud, Vu Minh Chien9, Terry Yue Zhuo10,11, Diganta Misra12,13, Ben Bogin14, Xuan-Son Vu15,16,17, Marzena Karpinska18, Arnav Varma Dantuluri, Wojciech Kusa33, Tommaso Furlanello, Rio Yokota1, Niklas Muennighoff, Suhas Pai19, Tosin Adewumi20, Veronika Laippala, Xiaozhe Yao21, Adalberto Junior, Alpay Ariyak22,23, Aleksandr Drozd24, Jordan Clive25, Kshitij Gupta12, Liangyu Chen, Qi Sun1, Ken Tsui, Noah Persaud, Nour Fahmy, Tianlong Chen8, Mohit Bansal8, Nicolò Monti26, Tai Dang18, Ziyang Luo27, Tien-Tung Bui28, Roberto Navigli3, Virendra Mehta29, Matthew Blumberg#30, Victor May#31,32, Huu Nguyen#32, Sampo Pyysalo#34 1Institute of Science Tokyo, 2MIT-IBM Watson Lab, 3Sapienza University of Rome, 4Babelscape, 5LAION, 6TU Darmstadt, 7hessian.AI, 8UNC Chapel-Hill 9Detomo Inc., 10CSIRO’s Data61, 11Monash University, 12Mila - Quebec AI Institute 13Carnegie Mellon University, 14Allen Institute for AI, 15DeepTensor AB, 16WASP Media & Language, 17Umeå University, 18University of Massachusetts Amherst, 19Hudson Labs, 20Luleå University of Technology, 21ETH Zurich, 22RunPod, 23OpenChat, 24RIKEN CCS, 25Chattermill AI, 26ASC27, 27Hong Kong Baptist University, 28DopikAI JSC 29University of Trento, 30GridRepublic, 31Chegg, 32Ontocord.AI, 33TU Wien, 34University of Turku taishi@rio.scrc.iir.isct.ac.jp, mayank.mishra2@ibm.com, tedeschi@diag.uniroma1.it, praty@cs.unc.edu diganta.misra@mila.quebec, mayvic@gmail.com, s.vu@deeptensor.ai, huu@ontocord.ai, sampo.pyysalo@utu.fi Equal contribution  #Equal mentoring
@@ 1 Introduction
Large Language Models (LLMs) are fundamental tools in artificial intelligence, powering applications such as machine translation, text summarization, dialogue systems, and code generation. These LLMs are pre-trained on extensive text data to enhance downstream task-specific adaptation. However, the excessive computational expense of pretraining LLMs creates barriers to access, constraining wider development.
Open-source initiatives such as Bloom (Scao et al., 2023), StarCoder (Li et al., 2023a), StarCoder-2 (Lozhkov et al., 2024), Pythia (Biderman et al., 2023), and OLMo (Groeneveld et al., 2024; Soldaini et al., 2024) have emerged to democratize access to pre-trained LLMs. These initiatives stimulate innovation, allowing researchers and developers to leverage existing advancements. However, despite their contributions, several significant challenges persist in the domain of open-source LLM development.
Primarily, several studies (Bang et al., 2023; Jiao et al., 2023; Hendy et al., 2023; Huang et al., 2023) have underscored the ongoing struggle of LLMs with non-English texts, particularly in low- or extremely low-resource languages. Given that the training data predominantly consists of English, as noted for instance by Brown et al. (2020) who reported that English accounts for 93% of GPT-3’s training corpus, there is a pressing need to promote the development of multilingual models to democratize LLMs and alleviate performance disparities across different languages (Chai et al., 2023). Secondly, continual pretraining – a technique involving further updating pretrained models on new data distributions to enhance their capabilities (Gupta et al., 2023; Fujii et al., 2024) – poses a significant challenge. While this approach could potentially enable life-long learning of large language models, it often leads to catastrophic forgetting, where the model loses previously acquired knowledge. This challenge is exacerbated when considering the continual pretraining of models across a diverse array of grammatical and lexical structures. Lastly, ensuring compliance with recent regulations mandating safe and secure AI development practices represents another critical aspect often overlooked in open-source LLM development, specifically, for multilingual models.
This paper presents Aurora-M, a novel open-source multilingual Large Language Model (LLM) with 15 billion parameters, tailored to address the aforementioned limitations. Aurora-M is designed to cater to five linguistically diverse languages: English, Finnish, Hindi, Japanese, Vietnamese, with a mix of code data. Aurora-M is continually pretrained from the StarCoderPlus model (Li et al., 2023a) on an extensive dataset comprising 435 billion tokens, resulting in a total training token count of an impressive 2 trillion tokens. This rigorous pretraining regimen equips Aurora-M with a comprehensive understanding of diverse languages and code. Moreover, safety is a fundamental design principle of Aurora-M. It stands out as the first open-source multilingual LLM fine-tuned on a comprehensive collection of human-reviewed safety instructions addressing concerns in the Biden-Harris Executive Order on Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence (WhiteHouse, 2023). This fine-tuning process not only addresses conventional red-teaming concerns (Ganguli et al., 2022; Perez et al., 2022) aimed at testing system vulnerabilities, but also aligns with the specific safety and security guidelines outlined in the Order.
@@
Our contributions can be summarized as follows.
@@
  • We construct a new red-teaming dataset, named “The Biden-Harris Redteam Dataset,” tailored to address concerns outlined in the Executive Order along with typical safety concerns. We then fine-tune Aurora-M on this dataset and evaluate on several safety benchmarks.
@@ 2 Datasets
Data Curation.
The continual pretraining process for training Aurora-M followed a carefully designed two-stage curriculum, as shown in Fig. 2. In the first stage, termed as Continual Auxiliary Pretraining (CAP), a large corpus of general multilingual web data was used to expose the model to diverse data, laying a robust foundation for subsequent training. The second stage, termed as Continual Alignment Tuning (CAT) employed a strategic data-mixing approach to bolster the model’s performance in targeted areas and align it with our predefined objectives. Following Taylor et al. (2022) and Li et al. (2023b), we also included publicly available instruction tuning datasets in both stages of training.
In CAP, we incorporated 377B tokens of processed and filtered web data from various sources, including Stack (Kocetkov et al., 2022), RefinedWeb (Penedo et al., 2023), RedPajama (Together, 2023), and a subset of the Pile (Gao et al., 2020). Additionally, multilingual data from HPLT (de Gibert et al., 2024), MC4 (Zhu et al., 2023a), Paracrawl (Ghussin et al., 2023), OSCAR (Abadji et al., 2022), along with Wikipedia (Foundation, 2023), and instruction tuning data from sources such as OpenAssistant (Köpf et al., 2023), APIBench (Patil et al., 2023), and OIG (LAION, 2023) were included.
For CAT, we opted for a greater percentage of code and a changed mix of high-quality public instruction datasets (Mishra et al., 2022a; Ding et al., 2023; Ivison et al., 2023), encompassing coding (Luo et al., 2023; Mishra et al., 2023a) and mathematical reasoning (Yu et al., 2023; Mishra et al., 2023b). The intention was to not overfit to the high quality instruction data, and thus the high quality data was used in CAT only. We also subsampled data from CAP for quality, as described below. Furthermore, we introduced a new safety instruction dataset named Biden-Harris Redteam, detailed in Section 4. The total dataset size for CAT is 58B tokens. We refer the reader to Fig. 2 for the distribution of languages in both training stages. The complete list of datasets is available in Appendix B.
@@
To remove toxic content and low-quality text, we applied filters similar to those used in Nguyen et al. (2023c) and Scao et al. (2023), such as stop-word proportions and text length. For all web text, we followed a process akin to Penedo et al. (2023) to remove low-quality content, including duplicate headers and footers. Additionally, in the CAT dataset, we further filtered web text with high proportions of symbols and numbers. In the case of RefinedWeb (Penedo et al., 2023), we utilized the RedPajama (Together, 2023) fastText classifier to retain English webpages resembling "high-quality" content similar to Wikipedia-linked articles. We trained and employed a similar classifier to filter other languages in our dataset, except for Finnish, where the procedure caused over-filtering, resulting in an excessively low sample volume post-filtering. To further enhance the quality of the RefinedWeb data, we adopted an approach detailed in Rönnqvist et al. (2021). We trained a fastText classifier [^1] Similar to https://github.com/TurkuNLP/register-labeling?tab=readme-ov-file and selectively subsampled web pages with over-represented registers, aiming to retain more "rare" text (e.g., lyrical or poetic text). This filtering process was specifically applied to English text due to the prohibitive slowness of our multilingual classifiers. Addressing this limitation represents an area for future research.
@@ 3 Model Training
For the training of Aurora-M, we maintained a consistent batch size of 2048 and a sequence length of 2048 tokens. The learning rate was linearly warmed up to $10^{-4}$ over 2,000 steps, followed by a cosine decay scheduler set to decay the learning rate to $10^{-5}$ by 120,000 steps. while optimization utilized the AdamW optimizer (Kingma and Ba, 2017; Loshchilov and Hutter, 2019) with coefficients $\beta_{1}=0.9$ and $\beta_{2}=0.95$. Additionally, Megatron-LM’s distributed optimizer with mixed precision training (Micikevicius et al., 2018) was used. Further training details can be found in the Appendix A.
@@ 4 Safety
LLMs can propagate harmful content, reinforce biases, or amplify misinformation. While users are responsible for assessing the potential risks of generated content, developers must prioritize legal and safety considerations, strengthening models against attacks that may bypass safety protocols.
In line with the Biden-Harris US Executive Order on AI (WhiteHouse, 2023), we curated the Biden-Harris Redteam Dataset, consisting of 5000 instruction-response pairs, addressing key concerns such as harm, cyber-attacks, CNBR risks, illegal acts, and privacy infringement. This dataset was created using a combination of filtering human preference data on harmlessness and template-based methods, with responses reviewed and edited for quality and safety. We used this dataset to instruction-tune Aurora-M and evaluated its safety levels before and after tuning. Details are provided in Section 5, with further dataset insights in Appendix C.
@@ 5.1 Evaluation Setup
We evaluated models across several English, Japanese, Finnish, Hindi, Vietnamese, and code-related benchmarks. For English, we used the Language Model Evaluation Harness (Gao et al., 2022) to assess tasks like OpenBookQA, TriviaQA, HellaSwag, SQuAD2.0, XWINO, and GSM8K. For Japanese, we followed swallow-llama and used llm-jp-eval (Han et al., 2024), covering JCommonsenseQA, JEMHopQA, and JSQuAD, among others. Finnish evaluation followed the method used in FinGPT with FIN-bench (Luukkonen et al., 2023a). We also evaluated Hindi and Vietnamese using the mlmm evaluation suite on tasks like HellaSwag and MMLU. For code evaluation, we utilized MBPP, HumanEval, MultiPL-E, and HumanEvalFix, and for safety, we employed datasets like the Biden-Harris Redteam Testset and DangerousQA. Detailed dataset descriptions and their corresponding evaluation metrics are provided in Appendix D.